A sample of nitrogen gas at 23°C and 785 mmHg occupies 3.84 mL. What is the volume, in mL, of the gas at STP?
1
Expert's answer
2020-04-03T11:50:59-0400

Solution:


"\frac {P_1 V_1} {T_1} = \frac {P_2 V_2} {T_2}"


"V_2=\frac {P_1 V_1 T_2} {P_2 T_1}"


"V_2=\frac {785 (mmHg) \times 3.84 (mL) \times 273 (K)} {760 (mmHg) \times 298 (K)}"


"V_2= 3.63 (mL)"


Answer:3.63 mL